N2 Use a triple integral to find the volume of the region in the first octant beneath the plane 40 pts 3x + y + 2z = 6
Answer #1

Na - The gives equation of the plane is 3x + y + 2z = 6 From the given equation the limits are the following os x=2, osy £ 6-= (36-367 +922) deze = 12 (4-4*+22) de Integrate with respect to x, we get = [ 48 – 42 + 2372 = 3 [4.) – 2.(4) + $ = (8-8 + $